Jason bought a home in Arlington, Texas, for $127,000. He put down 25% and obtained a...

  1. Jason bought a home in Arlington, Texas, for $127,000. He put down 25% and

obtained a mortgage for 30 years at 6%.

a.

What is Jason’s monthly payment? (Do not round intermediate calculations.

Round your answer to the nearest cent.)

  Monthly payment

$   

b.

What is the total interest cost of the loan? (Use the amortization worksheet on the financial calculator.)
Answer #1

Cost of Home = $127,000

Down Payment = 0.25(127,000) = $31,750

Loan Amount = 127,000 - 31,750

Loan Amount = $95,250

Time Period = 30 years

Interest Rate = 6%

Calculating Monthly Payment on Loan,

Using TVM Calculation,

PMT = [PV = 95,250, T = 360, FV = 0, I = 0.06/12]

PMT = $571.07

Monthly Payment = $571.07

Total Interest Cost = Total Amount Paid - Loan Amount

Total Interest Cost = 360(571.07) - 95,250

Total Interest Cost = $110,335.20